Question 1) Find I = z +2 3z - 2 + 3i 22 + (2i - 2)2 - 4i ] dz, C:\z| = 3, CW

a. 4πί b. 8πί C. 2πί d. -2π(3 +i) e. 0.0 f. ο g. -4πί h. 6π
Answer #1

Date Page NO I-61 C:111=3 , Cw J.L2+2 + 32-2+3; de z²+(21-2)=-40 are af z=-2, - bon polon S x²+(21-2) Z-41=0 7²22+Riz-yito 2=So I = I + In I = -2Ti - bril I - - 8ril AnswerThe answer must be -8πi because on changing the direction of integral i.e. clockwise there must be a negative sign. In standard notation we consider direction of closed loop to be counter-clockwise direction (anticlockwise) as positive.

There might be misprint in the options.